# Improving LLM Personas via Rationalization with Psychological Scaffolds

### v1

@@ C. Refused to answer
Table 2: Snippets of rationales generated by GPT-4 for a career satisfaction judgment according to the scaffolds used in PB&J. Full examples are shown in Appendix G, along with user demographics
@@
Our persona description takes as input a set of seed judgments, for which we generate rationales. For OpinionQA, these seed judgments are a held-out set of questions and a users’ set of corresponding answers to those questions. The questions and answers are then converted into declarative forms (e.g., If a user responds No to “Are you currently focused on your professional life and career?”, the declarative form would be I’m not very focused on my professional life right now). For MovieLens, we use movie ratings provided by the user with earlier timestamps for predicting ratings by the same user with later timestamps. Each judgment consists of a movie, its rating, and a short description of the movie in terms of its plot synopsis, actors, and directors. For our main experiments, a fixed set of 8 seed judgments are provided in the persona description.
@@
We use dataset opinion questions and movies that are not a part of the seed judgments to evaluate the personas. Since our personas are customized to each real user, we calculate accuracies and standard deviation for both OpinionQA and MovieLens, macro-averaged for each user.
Models.
We use two models of varying sizes — GPT-4  [^2] Last accessed on 2 December 2024, GPT-4 points to gpt-4-0613.OpenAI et al. (2024) and Mistral 0.2 Instruct 7B Jiang et al. (2023). All results displayed are based on prompting these models, without any fine-tuning.
Baselines.
Following prior work, we experiment with different variants of building persona descriptions as baselines, which are mostly ablations of PB&J. Santurkar et al. (2023) propose two methods – one where LLMs are prompted without any user information (No Persona) and with demographics of the user (Only Demographics). Hwang et al. (2023) add judgments to the persona descriptions without (Only Judgments) and with demographics (Demographics + Judgments). For the latter setting, we include results using Chain-of-thought reasoning Wei et al. (2023) as $\textsc{Demographics + Judgments}_{CoT}$.
@@
Existing persona LLM baselines are (almost) equivalent.
Table 3 presents macro-averaged accuracies for various baseline methods across models and datasets. Without any persona information (No Persona condition), both GPT-4 and Mistral 7B perform significantly worse, highlighting the necessity of incorporating some form of user context. The performance drop is particularly stark for GPT-4, likely due to its safety guardrails, which often lead to abstentions from expressing opinions or preferences Chen et al. (2023). This is evidenced by frequent responses such as, "This is a subjective question and the answer will vary" or "As an AI, I don’t have personal experiences", which contribute to its lower accuracy. [^3] This behavior does not occur for all questions but has a substantial impact on overall performance. Among other baselines, demographic attributes (Only Demographics) generally provide better performance than judgments alone (Only Judgments), especially for GPT-4. The best-performing baseline varies: While Demographics + Judgments is the strongest baseline on OpinionQA, its advantage does not generalize to MovieLens, where Only Demographics performs comparably or better. This suggests that the utility of judgments depends on the nature of the task—while they offer some alignment in opinion-based questions, they are less effective in preference-based settings like MovieLens.
PB&J provides significant improvements in aligning personas.
Across both datasets (Table 3), PB&J consistently outperforms baselines, demonstrating the effectiveness of rationales. We show an example generation of rationales in Table 2. On OpinionQA, rationale scaffolds yield 9.7%–10.8% improvements over Demographics + Judgments, with Primal World Beliefs achieving the highest accuracy, closely followed by Experiences. Even unstructured rationales (No Scaffold) surpass Demographics + Judgments, reinforcing the value of rationales in persona alignment. Trends hold across models, with GPT-4 maintaining a stronger baseline.
In MovieLens, psychological scaffolds are even more critical. While GPT-4 performs better overall, Mistral 7B benefits substantially from Schwartz Theory of Basic Human Values and Primal World Beliefs, which provide the largest gains. Unlike in OpinionQA, No Scaffold fails to improve over Demographics + Judgments, indicating that psychological scaffolds may be particularly useful for modeling subjective preferences. Notably, Primal World Beliefs consistently ranks among the top scaffolds across datasets, suggesting that belief-driven rationalization is a powerful tool for persona modeling. However, Schwartz Theory of Basic Human Values excels in MovieLens, highlighting that different tasks may benefit from different psychological frameworks.
5 Using PB&J with human-written rationales
@@
Table 4: Incorporating human-written rationales in PB&J: Human-written rationales for OpinionQA judgments consistently perform baselines and all LLM-generated rationales except Primal World Beliefs.
Table 3 demonstrates that LLM-generated rationales effectively improve persona alignment in opinion prediction and movie preference tasks. However, we hypothesize that the observed improvements stem not from the LLM-generated rationales themselves, but from the additional context they provide, enabling more accurate generalization for each user. To test this, we conducted a pilot study using a subset of 30 questions from OpinionQA. We recruited 100 users, collected their demographic information and responses, and asked them to provide rationales for their first nine answers. These rationales were open-ended, allowing participants to explain their opinions based on personal experiences, beliefs, or personality traits without any imposed structure. The remaining 21 questions were used for evaluation. More details about this pilot are discussed in Section B.3
Table 4 reports results on this subset comparing baseline methods, LLM-generated rationales, and Human Written rationales using GPT-4. Human-written rationales outperform all baselines and PB&J variants, with the exception of Primal World Beliefs, where the difference is not statistically significant. This highlights the effectiveness of human-generated explanations in improving persona alignment, even with limited judgments. Additionally, the comparable performance of Primal World Beliefs rationales to human-written ones suggests that carefully designed scaffolds can help LLM-generated rationales approximate human rationale quality. These findings emphasize the role of rationalization in improving persona LLMs, whether through human input or well-constructed LLM scaffolds.
@@ 6 Discussion
LLM-generated rationales help with a limited budget of judgments.
Unlike demographics, user judgments take time to collect, making it crucial to assess persona alignment with limited data (Section 4). As shown in Figure 3, PB&J with Primal World Beliefs outperforms both baselines (Only Judgments and Demographics + Judgments) across all judgment counts. Even with just two judgments, PB&J surpasses baselines far more, highlighting the value of LLM-generated rationales in low-data settings. As the number of judgments increases, the performance of all methods improves. However, the rate of improvement for all diminishes after 8 judgments. This indicates that while all methods benefit from additional user judgments, PB&J maximizes its potential gains earlier due to the contextual richness provided by LLM-generated rationales.
@@
A robust personalization method should improve performance across diverse user groups rather than relying on gains from specific demographics. To assess this, we compare PB&J (with Primal World Beliefs scaffolds) to Demographics + Judgments across education, race, income, and gender in Figure 4 (see full results in Figure 5 in the Appendix). PB&J consistently outperforms the baseline across all demographics. While gains are notable for users with postgraduate education and those identifying as Asian or Mixed Race, improvements extend across all education levels, racial groups, income brackets, and genders. By incorporating structured rationales, PB&J improves with diverse user perspectives, demonstrating broad effectiveness without demographic biases.
@@
In cases where user demographics are unavailable or where the risks of model bias outweigh the utility of demographic information, PB&J can improve performance of personalization from judgments alone. Rationalizing Only Judgments with Primal World Beliefs results in an 11.64% absolute improvement on OpinionQA and 1.1% on MovieLens (GPT-4). While still under-performing demographic-based methods, this result indicates that deeper reasoning about user behavior alone may obviate the need for coarse and potentially biasing demographic information.
@@ 7 Related Work
Personalizing LLMs.
Recent works have used LLM personas to simulate user behavior in psycholinguistic experiments Aher et al. (2023); Karra et al. (2022b), or other social science experiments Filippas et al. (2023); Argyle et al. (2023). Specifically, the use of LLMs to simulate user responses to surveys and other value-laden scenarios, using existing user information like demographic background has been gaining increasing attention Santurkar et al. (2023); Hwang et al. (2023); Durmus et al. (2024); Chuang et al. (2024); Do et al. (2024); Sun et al. (2025); Moore et al. (2024); Dillion et al. (2023); Tjuatja et al. (2023); Balepur et al. (2025) Recently, that attention has shifted more towards synthetically augmenting persona information Moon et al. (2024); Simmons (2022) or completely synthesizing personas generated from seed human data Park et al. (2022, 2024, 2023); Ge et al. (2024).
@@
Psychologists have long investigated how different sociological and psychological aspects influence personality-driven narratives of an individual McAdams (1993); Bruner (1991); Pennebaker and King (1999). Recently, researchers have focused on eliciting psychological markers for evaluating LLMs Hilliard et al. (2024); Karra et al. (2022a); Serapio-García et al. (2023). Very few works actually incorporate these principles in an LLM persona itself Moon et al. (2024); Park et al. (2024).
@@
Previous works have focused on generating reasoning chains or rationales by either prompting language models Wei et al. (2023); Saha et al. (2023) or learning to generate rationales by fine-tuning on such data Wiegreffe et al. (2022); Ramnath et al. (2024). Recently, there has been a surge in built-in reasoning capabilities in language models via extensive ‘learning-to-reason’ steps Zelikman et al. (2022); DeepSeek-AI et al. (2025). It is important to note that these prior works focus on reasoning about the data, by primarily asking - Why did the model produce a certain output? Instead, PB&J can be seen as reasoning about people rather than the data, and scaffolds are a way to instruct the model how to reason.
@@ 8 Conclusion
This work introduces PB&J, a framework that improves LLM personas by incorporating psychologically-inspired rationales to explain user judgments. By leveraging structured psychological scaffolds, PB&J improves LLM persona performance across diverse opinion prediction and preference modeling tasks and also at par with human-written rationales. Additionally, PB&J remains effective even with limited user judgments, highlighting its potential for real-world personalization applications, with limited user history.
@@
While PB&J improves personas through LLM-generated rationales, it relies solely on zero-shot prompting for both rationale generation and downstream predictions. While this allows for flexibility and adaptability across users, it may not fully capture the complexity and depth of individual reasoning. Fine-tuning LLMs on human-written rationales could further improve personalization by enabling models to learn user-specific patterns rather than relying solely on generated rationales. Additionally, while these rationales improve performance, we cannot validate their fidelity to actual user reasoning, as no ground-truth rationales are available. This limitation is inherent to synthetic-data based persona modeling Moon et al. (2024); Park et al. (2024), where plausible explanations generated by LLMs may align well with observed user behavior but not necessarily reflect the true underlying motivations.
@@
Our study primarily evaluates PB&J on U.S.-based user populations, as both OpinionQA and our human pilot study consist of participants located in the United States. Our study was conducted under the guidance of an ethics review board. Additionally, the subset of users selected from MovieLens also resides in the U.S. While this ensures consistency in evaluation, it limits the generalizability of our findings to more diverse global populations. Future work should explore the effectiveness of PB&J across different cultural and linguistic contexts to ensure broader applicability. Since our datasets involve personal judgments on opinion-based questions, some generated responses may reflect viewpoints that could be offensive or controversial. While we do not directly intervene in the LLMs’ generation of rationales, it is crucial to recognize that models can inherit biases present in both training data and user-generated inputs. Finally, as with any system that models human behavior, there are concerns around user privacy. While our work does not use real user data beyond voluntary survey responses, deploying such approaches in real-world settings would require careful consideration of data collection practices, consent mechanisms, and safeguards against potential misuse.
Appendix A PB&J prompts

### v2

@@ 4 Experiments
Table 2: Snippets of rationales generated by GPT-4 for a career satisfaction judgment according to the scaffolds used in PB&J. Full examples are shown in Appendix D, along with user demographics
Task and Datasets.
We conduct experiments on two datasets - OpinionQA Santurkar et al. (2023) and MovieLens Harper and Konstan. (2015). OpinionQA is a collection of American public opinion surveys Pew Research Center (2025) conducted by PEW Research containing user-selected answers to multiple-choice questions in 15 different topics ranging from food safety to guns, with rich accompanying demographic attributes. The task is to predict a user’s answer for a question. For OpinionQA, we use a subset of 750 users, and 10 test questions for each user. MovieLens contains timestamped movie ratings (between 1-5) corresponding to individual users, with gender, age, occupation and location of each user. The task is to predict a user’s rating for a movie. We use a subset of 100 users and 10 test movies for each user. Further details and splits of the dataset are presented in Appendix C. Example inputs-outputs of each of these datasets are shown in Figure 2.
@@
Our persona description generates rationales for a set of seed judgments. For OpinionQA, these seed judgments are a ‘train’ set of questions and users’ corresponding answers to them. The questions and answers are then converted into declarative forms (e.g., If a user responds No, not really to “Are you currently focused on your professional life and career?”, the declarative form would be I’m not very focused on my professional life right now). For MovieLens, we use movie ratings provided by the user with earlier timestamps for predicting ratings by the same user with later timestamps. Each judgment consists of a movie, its rating, and a short description of the movie consisting of its plot synopsis, actors, and directors. For our main experiments, a fixed set of 8 seed judgments are provided in the persona description.
@@
We use a ‘test’ set of opinion questions and movies to evaluate the personas. Since our personas are customized to each real user, we calculate accuracies and standard deviation for both OpinionQA and MovieLens, macro-averaged for each user.
@@
Language Models.
We use two LMs of varying sizes — GPT-4  [^1] Last accessed on 2 December 2024, GPT-4 points to gpt-4-0613.OpenAI et al. (2024) and Mistral 0.2 Instruct 7B Jiang et al. (2023). All results displayed are based on prompting these LMs, without any fine-tuning.
Baselines.
Following prior work, we experiment with different variants of building persona descriptions as baselines, Santurkar et al. (2023) propose two methods – one where LMs are prompted without any user information (No Persona) and with demographics (Only Demographics). Hwang et al. (2023) add judgments to the persona descriptions without (Only Judgments) and with demographics (Demographics + Judgments). For the latter setting, we also include results using Chain-of-thought reasoning Wei et al. (2023) as $\textsc{Demographics + Judgments}_{CoT}$, where $LM_{\mathcal{T}}$ is prompted to reason about the test question using the persona information in order to arrive at the answer. This is $LM_{\mathcal{T}}$’s default rationale that is not explicitly grounded in any psychological scaffold.
Existing persona LM baselines are (almost) equivalent.
Table 3 presents macro-averaged accuracies for various baseline methods across models and datasets. Without any persona information (No Persona condition), both GPT-4 and Mistral 7B perform significantly worse, highlighting the necessity of incorporating some form of user context. The performance drop is particularly stark for GPT-4; due to its safety guardrails, GPT-4 abstains frequently from expressing opinions and preferences Chen et al. (2023). This leads to responses such as "This is a subjective question and the answer will vary" or "As an AI, I don’t have personal experiences" [^2] Such behavior does not occur for all questions but has a substantial impact on overall performance.. Demographic attributes (Only Demographics) are generally more predictive than judgments alone (Only Judgments) with Demographics + Judgments as the best performing baseline, especially for GPT-4. However, it is interesting to see that providing additional tokens to reason via Chain-of-Thought rationales in $\textsc{Demographics + Judgments}_{CoT}$ also yields no significant improvements over Demographics + Judgments.
@@
Table 4: Incorporating human-written rationales in PB&J: Human-written rationales for OpinionQA judgments consistently outperform baselines and all LM-generated rationales except Primal World Beliefs.
PB&J provides significant improvements for personas.
Across both datasets (Table 3), PB&J consistently outperforms the baselines, demonstrating the effectiveness of plausible, synthetic rationales (example generations in Table 2). On OpinionQA, scaffolded rationales yield significant improvements over Demographics + Judgments, even after generating Chain-of-Thought rationales, with Primal World Beliefs achieving the highest accuracy, closely followed by Experiences. Even unstructured rationales (No Scaffold) surpass $\textsc{Demographics + Judgments}_{CoT}$. These trends hold across LMs.
For MovieLens, psychological scaffolds are even more critical. While GPT-4 performs better overall, smaller LMs like Mistral 7B benefit substantially from Schwartz Theory of Basic Human Values and Primal World Beliefs scaffolds. Primal World Beliefs consistently ranks among the top scaffolds across datasets, however, Schwartz Theory of Basic Human Values excels in MovieLens, highlighting that different tasks may benefit from different psychological frameworks. Lastly, we also experimented with combinations of scaffolds, including synthesizing summary rationales from all scaffolds, but these were less performant (Appendix F).
5 Using PB&J with human-written rationales
Table 3 demonstrates that plausible yet synthesized LM rationales effectively improve user personas. However, we hypothesize that the observed improvements stem not from the LM-generated rationales themselves, but from the additional context they provide, enabling more accurate generalization for each user. To test this, we conducted a pilot study using a subset of 30 questions from OpinionQA. We recruited 100 users to respond to the subset of questions, collected their demographics and asked them to provide rationales for the first nine responses. These rationales were unconstrained, allowing participants to explain their judgments based on personal experiences, beliefs, or personality traits without any imposed structure. The remaining 21 questions were used for evaluation. Details are in Section C.3.
Table 4 reports results on this subset, with the inclusion of Human Written using GPT-4. Human Written rationales outperform all baselines and PB&J variants, with the exception of Primal World Beliefs, where the difference is not statistically significant. This highlights that while PB&J generates synthetic rationales, plausible and carefully selected scaffolds hold similar predictive power to Human Written rationales.
@@ 6 Discussion
LM-generated rationales help even with a limited budget of judgments.
Unlike demographics, user judgments take time to collect, making it crucial to assess personas with limited data (Section 4). As shown in Figure 3, PB&J with Primal World Beliefs outperforms both baselines (Only Judgments and Demographics + Judgments) across varying # of judgments. Even with just two judgments, PB&J surpasses baselines far more, highlighting the value of LM-generated rationales in low-data settings. As the number of judgments increases, the performance of all methods improves. However, the rate of improvement for all diminishes after 8 judgments. This indicates that while all methods benefit from additional user judgments, PB&J maximizes its potential gains earlier due to the contextual richness provided by LM-generated rationales.
@@
A robust personalization method should improve performance across diverse user groups rather than relying on gains from specific demographics. To assess this, we compare PB&J (with Primal World Beliefs scaffolds) to Demographics + Judgments across education, race, income, and gender in Figure 4 (splits across all demographics in Figure 5). PB&J consistently outperforms the baseline across all demographics. While gains are notable for users with postgraduate education and those identifying as Asian or Mixed Race, improvements extend across all education levels, racial groups, income brackets, and genders. By incorporating scaffolded rationales, PB&J improves with diverse user perspectives, demonstrating broad effectiveness with synthetic rationales.
@@
In cases where user demographics are unavailable or where the risks of bias outweigh the utility of demographic information, PB&J can improve performance of personalization from judgments alone. Rationalizing Only Judgments with Primal World Beliefs results in an 11.64% absolute improvement on OpinionQA and 1.1% on MovieLens (GPT-4). While still under-performing demographic-based methods, this result indicates that deeper reasoning about user behavior alone may obviate the need for coarse and potentially biasing demographic information.
Effect of reasoning length on performance.
[…]
@@ 7 Related Work
Personalizing LM.
Recent works have used LM personas to simulate behavior in psycholinguistic and other social science experiments Aher et al. (2023); Karra et al. (2022b); Filippas et al. (2023); Argyle et al. (2023). Specifically, the use of LMs to simulate user responses to surveys, using existing user information like demographic background has been gaining increasing attention Santurkar et al. (2023); Hwang et al. (2023); Durmus et al. (2024); Chuang et al. (2024); Do et al. (2024); Sun et al. (2025); Moore et al. (2024); Dillion et al. (2023); Tjuatja et al. (2023); Balepur et al. (2025) Recently, that attention has shifted more towards synthetically augmenting persona information Moon et al. (2024); Simmons (2022) or completely synthesizing personas generated from seed human data Park et al. (2022, 2024, 2023); Ge et al. (2024).
@@
Psychologists have long investigated how different sociological and psychological aspects influence personality McAdams (1993); Bruner (1991); Pennebaker and King (1999). Recently, researchers have focused on eliciting psychological markers for evaluating LMs Hilliard et al. (2024); Karra et al. (2022a); Serapio-García et al. (2023). Very few works actually incorporate these principles in an LM persona itself Moon et al. (2024); Park et al. (2024).
@@
Previous works have focused on generating reasoning chains or rationales by either prompting language models Wei et al. (2023); Saha et al. (2023) or learning to generate rationales by fine-tuning on such data Wiegreffe et al. (2022); Ramnath et al. (2024). Recently, there has been a surge in built-in reasoning capabilities in language models via extensive steps Zelikman et al. (2022); DeepSeek-AI et al. (2025).
@@ 8 Conclusion
This work introduces PB&J, a framework that improves LM personas by incorporating plausible, yet synthetic rationales to explain user judgments. By leveraging psychological scaffolds, PB&J improves LM persona accuracy across diverse opinion prediction and preference modeling tasks, while performing at best close to human-written rationales. Additionally, PB&J remains effective even with limited user judgments, highlighting its potential for real-world personalization applications, with limited user history.
@@
While PB&J improves personas through plausible yet synthetic LM-generated rationales, it relies solely on zero-shot prompting for both rationale generation and downstream predictions. While this allows for flexibility and adaptability across users, it may not fully capture the complexity and depth of individual reasoning. Fine-tuning LMs on human-written rationales could further improve personalization by enabling models to learn user-specific patterns rather than relying solely on generated rationales. Additionally, while these rationales improve performance, we cannot validate their fidelity to actual user reasoning, as no ground-truth rationales are available. This limitation is inherent to synthetic-data based persona modeling Moon et al. (2024); Park et al. (2024), where plausible explanations generated by LMs may align well with observed user behavior but not necessarily reflect the true underlying motivations. However, we emphasize and caution that synthesized rationales can be a plausible reason for a judgment, but may not be the exact reason used by a user. We provide further analyses about this in Appendix I.
@@
Our study primarily evaluates PB&J on U.S.-based user populations, as both OpinionQA and our human pilot study consist of participants located in the United States. Our study was conducted under the guidance of an ethics review board. Additionally, the subset of users selected from MovieLens also resides in the U.S. While this ensures consistency in evaluation, it limits the generalizability of our findings to more diverse global populations. Future work should explore the effectiveness of PB&J across different cultural and linguistic contexts to ensure broader applicability. Since our datasets involve personal judgments on opinion-based questions, some generated responses may reflect viewpoints that could be offensive or controversial. While we do not directly intervene in the LMs’ generation of rationales, it is crucial to recognize that models can inherit biases present in both training data and user-generated inputs. Finally, as with any system that models human behavior, there are concerns around user privacy. While our work does not use real user data beyond voluntary survey responses, deploying such approaches in real-world settings would require careful consideration of data collection practices, consent mechanisms, and safeguards against potential misuse.
Acknowledgments
We thank anonymous reviewers and lab members at Apple and USC NLP for their feedback on this work. This work was also supported in part by the National Science Foundation under grant IIS-2403437, the Simons Foundation, and the Allen Institute for AI. Any opinions, findings, conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. This work was partially done while B. Joshi was at Apple. and S. Swayamdipta was a visitor at the Simons Institute for the Theory of Computing. B. Joshi was also supported by the Apple Scholars in AL/ML PhD Fellowship.
Appendix A More on Psychological Scaffolds
A rich body of research has extended, debated, and validated each of the psychological scaffolds presented earlier. Among these, Primal World Beliefs stands out as the most linguistically motivated framework, which may explain its strong performance in our experiments. As described in Clifton et al. (2019), researchers analyzed historical texts and over 80K tweets using topic modeling and concept extraction to identify statements about how people perceive the world. These statements were then categorized through expert coding, consultation with social scientists, and discussions with religious focus groups, leading to the identification of 26 fundamental “primal world beliefs". These beliefs encapsulate deep-seated assumptions individuals hold about the world, such as whether the world is inherently safe or dangerous, simple or complex, and abundant or limited. Expanding on this work, Vu et al. (2022a) introduced a Latent Beliefs Model, which leverages transformer-based embeddings and a modified GPT-2 model to automatically infer latent dimensions of human beliefs from social media text. This data-driven discovery of worldviews underscores the linguistic basis of primal beliefs and their connection to naturally occurring human rationales.
Another well-established framework in social psychology is Schwartz Theory of Basic Human Values Schwartz (1992); Schwartz et al. (2012). This theory posits that human values, that are fundamental guiding principles, are organized along universal motivational dimensions that drive behavior. Schwartz identifies ten broad value categories, such as self-direction (independence of thought and action), benevolence (concern for others’ welfare), and power (desire for dominance or control). These values are structured in a circular model, where adjacent values are more compatible, and opposing values (e.g., security vs. stimulation) tend to be in tension. A key aspect of Schwartz’s values is their cross-cultural validation; extensive empirical studies have shown that these value dimensions hold across diverse populations, making them a robust framework for modeling user judgments in an LM setting. Unlike primal beliefs, which describe broad worldviews, Schwartz’s values provide a structured way to infer decision-making tendencies and moral considerations, making them particularly useful for understanding user preferences in ethical or societal questions.
Appendix B PB&J prompts
PB&J is primarily an inference-time prompt-based strategy to improve LM personas, without requiring any fine-tuning. We pick the basic structure of our prompt from Hwang et al. (2023). For all psychological scaffolds, we try multiple variations of prompts with varying levels of instructions, definitions and examples. Varying instructions for scaffolds does not lead to significant changes in performance. For example, adding more detailed instructions for Big 5 Personality Traits leads to a small, yet insignificant increase of 0.6 points for the OpinionQA Human Written subset.